OpenStudy (anonymous):

Which of the following is equivalent to a a^(1/3)*b^(3/4)*c^(5/2)? note: example: √7:( means 7th root √12:(ab^3c^3) 12(ab^3c^5) (abc)^(43/12) √9:(a^3b^4c^5) (c^2)√12:(a^4b^9c^6)

OpenStudy (anonymous):

Fractional exponents are power-over-root.

OpenStudy (anonymous):

ok, and?

OpenStudy (anonymous):

Just a sec, I'll type out the preliminary steps..

OpenStudy (anonymous):

\[a^{1/3}b^{3/4}c^{5/2} = a^{4/12}b^{9/12}c^{30/12} = \sqrt[12]{a^4b^9c^{30}}\]

OpenStudy (anonymous):

Can you find the answer from there?